OpenStudy (anonymous):

Solve 8t + 16 > 16t - (12t - 20). A. {t| t > 1} B. {t| t > -1} C. {t| t < 4} D. {t| t > 4}

OpenStudy (anonymous):

start from here... (-)*(12t-20)= ?

OpenStudy (anonymous):

can you do that?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

(-)*(-)=? minus times minus is equal to what? :)

OpenStudy (anonymous):

+

OpenStudy (anonymous):

ok how about (-)*(+)=?

OpenStudy (anonymous):

-

OpenStudy (anonymous):

ohh.. sorry I wasn't here.. yes that's right , so we have: 8t + 16 > 16t - (12t - 20). 8t + 16 > 16t - 12t + 20 8t + 16 > 4t + 20 now add -4t to both sides and -16 to both sides.. tell me what you get.